OpenStudy (anonymous):

find b if v=<1/5,b> is a unit vector

OpenStudy (zarkon):

find \[\|v\|\]

OpenStudy (anonymous):

What is v?

OpenStudy (anonymous):

v isn't given. that is all the question states

OpenStudy (anonymous):

how do i find IIvII if i don't know b?

OpenStudy (zarkon):

the same way you would do it if you did know b

OpenStudy (anonymous):

I agree ^-^ because it is asking the sam thing kinda

OpenStudy (zarkon):

\[\|<\alpha,\beta>\|=\sqrt{\alpha^2+\beta^2}\]

OpenStudy (anonymous):

ok and set that equal to zero and solve for b?

OpenStudy (zarkon):

no...set equal to 1...it is supposed to be a unit vector

OpenStudy (zarkon):

solve \[\|<1/5,b>\|=\sqrt{(1/5)^2+b^2}=1\] for b

OpenStudy (anonymous):

so would it be 4/5?

OpenStudy (zarkon):

no

OpenStudy (anonymous):

1/5 squared is 1/25

OpenStudy (anonymous):

square root of that is 1/5

OpenStudy (anonymous):

subratct that from 1

OpenStudy (zarkon):

\[\sqrt{(1/5)^2+b^2}=1\] \[(1/5)^2+b^2=1\] \[b^2=1-(1/5)^2=1-1/25=24/25\] \[b=\pm\sqrt{24/25}\]

myininaya (myininaya):

zarkon when you are done here can you see if you know this http://openstudy.com/groups/mathematics#/groups/mathematics/updates/4e6832ed0b8b4a2b95c7826f

myininaya (myininaya):

because i don't

OpenStudy (anonymous):

my math lab says that b is wrong?

myininaya (myininaya):

so what does math lab say?

OpenStudy (anonymous):

sorry it was right, i just typed it in wrong